# Spin-orbit misalignment and residual eccentricity are evidence that neutron star-black hole mergers form through triple star evolution

### v1

@@ 2.3 Gravitational three-body dynamics
Near a large eccentricity excitation of the inner binary the timescale for changes in the orbital elements could become shorter than the outer or inner orbital period, rendering the secular equations of motion inaccurate and likely underestimating the number of eccentric mergers by a few ten percent compared to direct three-body integration (e.g., F. Antonini et al., 2014; E. Grishin et al., 2018; C. Hamilton & R. R. Rafikov, 2024). Moreover, we do not attempt to model tidal deformation and mass-transfer of NSs which may alter the properties of NSBHs near the merger (M. B. Davies et al., 2005; F. Foucart et al., 2024; Y. Zenati et al., 2025), which is beyond the scope of this work.
2.4 Initial conditions
@@
On the zero-age-main-sequence (ZAMS) we sample the initial mass of the primary star in the inner binary from an initial mass function $p(m_{1})\propto m_{1}^{-2.3}$ between $10$ and $100\,\rm M_{\odot}$ (P. Kroupa, 2001). We follow the observationally inferred distribution of massive binary stars by H. Sana et al. (2012) and sample the inner binary mass ratio $q_{1}=m_{2}/m_{1}$ from a power-law $p(q_{1})\propto q_{1}^{-0.1}$ and its orbital period from $p(\log P_{1}/{\rm days})\propto(\log P_{1}/{\rm days})^{-0.55}$. We reject any inner binary which would overflow its Roche-lobe at ZAMS. We assume that the inner binaries are initially circular ($e_{1}=0$) and sample the outer binary from a thermal eccentricity distribution, a log-uniform semi-major axis distribution, and the outer binary mass ratio $q_{2}=m_{3}/m_{1}$ from a broken power-law $p(q_{2})\propto q_{2}^{-1.2}$ below $q_{2}<0.3$ and $\propto q_{2}^{-2.0}$ above $q_{2}<0.3$ (M. Moe & R. Di Stefano, 2017), limited to $0.08\,{\rm M_{\odot}}<m_{3}<5\,\rm{\rm M_{\odot}}$. For any given inner binary, we repeat sampling the outer binary until the triple is dynamically stable (R. A. Mardling & S. J. Aarseth, 2001). We then discard any triple with an outer semi-major axis above $a_{2}>5\times 10^{4}\,\rm AU$ because they are likely disrupted by stellar fly-bys and tides from the host galaxy (Y.-F. Jiang & S. Tremaine, 2010; K. El-Badry & H.-W. Rix, 2018; C. Hamilton & S. Modak, 2024; J. Stegmann et al., 2024). Finally, the orbital orientations of the inner and outer orbital frame are sampled from random isotropic distributions, leading to a random mutual inclination between the inner and outer orbit. Using the procedure described above, we construct a total number of $N_{\rm tot}=10^{7}$ stellar triples which we evolve in this work.
@@ 3 Results
Starting on the ZAMS, the orbits of stellar triples have to survive two consecutive SNe and any inner binary mass transfer episode in order to successfully develop stable NSBH triples that could eventually lead to tertiary-driven mergers. In Fig. 2 we show how the properties of surviving systems change across different times during the stellar evolution. We find that a few $\sim 0.1\,\%$ of all stellar triples initiated at ZAMS (grey patches) survive the stellar evolution and form a stable NSBH triple (red lines) and about $\sim 40\,\%$ thereof subsequently lead to a tertiary-driven merger (red patches). The latter fraction is similar to the merger fraction found by G. Fragione & A. Loeb (2019a) for their population of wider NSBH triples. During stellar evolution the NS formation at second SN is the most detrimental event to the triple stability and critically truncates the parameter distribution of the survivors. On the one hand, the natal kick of the newly formed NS poses the risk to disrupt the inner binary producing a dearth of stable NSBH inner binaries (blue line) above $a_{1}\gtrsim 10^{2}\,\rm AU$ (left panel). On the other hand, the mass-loss associated with the SN imparts a kick to the inner binary centre-of-mass which causes a steep decline in the distribution of $a_{2}/a_{1}$ of surviving NSBH triples above $\gtrsim 10^{2}$ (right panel). The stellar evolutionary constraints are further illustrated in Fig. 3 showing that the surviving NSBH triples occupy a somewhat narrow strip in $a_{2}$ versus $a_{1}$ which is predominantly determined by the mass-loss associated with the second SN and the need for the outer orbit to survive it. Consequently, we find different natal kick models only to affect the upper limit on the inner binary semi-major axis of surviving NSBH triples but not substantially changing the semi-major axis ratio distribution, as shown in Fig. 4.
@@
As a result, the surviving NSBH triples naturally populate a distinct region of parameter space that facilitates the formation of tertiary-driven mergers with spin-orbit misalignment and residual eccentricity. On the one hand, the SN kick imparted to the newly formed NS disrupts triples in wide architectures and is only withstood by triples with relatively small inner and outer semi-major axes. One the other hand, NSBH triples are characterised by a small inner binary mass ratio $q=m_{\rm NS}/m_{\rm BH}\ll 1$ since $m_{\rm NS}\ll m_{\rm BH}$. Therefore, we obtain tertiary-driven NSBH mergers that are, by and large, characterised by (i) strong octupole effects, (ii) a non-adiabatic BH spin evolution, (iii) a peculiar dynamics in which the NSBH mergers “plunge” (C. L. Rodriguez & F. Antonini, 2018) directly into a highly eccentric, GW-dominated phase without being significantly quenched by apsidal Schwarzschild precession of the inner binary, and (iv) short delay times from ZAMS to merger. Moreover, (v) the evolution of the NSBH progenitor stars is dominated by close stellar binary interactions, as explained below.
@@
    Octupole effects enable mergers across a wide range of initial relative inclinations: The secular equations of motion of hierarchical triples are obtained by expanding the three-body Hamiltonian in semi-major axis ratio $a_{1}/a_{2}$ (e.g., R. S. Harrington, 1968). To the lowest (“quadrupole”) order large inner binary eccentricities are only excited if the relative inclination between the inner and outer orbital planes is near-perpendicular. Large eccentricities across a much wider range of relative inclinations can be excited if the next-order (“octupole”) terms becomes relevant (e.g., E. B. Ford et al., 2000; B. Liu et al., 2015; S. Naoz, 2016) whose relative strength can be characterised as (S. Naoz et al., 2013)
@@
    As shown in Fig. 5, $\epsilon_{\rm oct}$ of the surviving triples can become as large as $\epsilon_{\rm oct}\lesssim 0.1$ (left panel) because the inner binary masses of NSBHs are very unequal, the semi-major axis ratio of surviving NSBH triples is relatively compact, and the thermal eccentricity distribution of the outer binary disfavours circular outer orbits (middle panel). NSBH mergers predominantly occur with nearby tertiaries where $a_{2}(1-e_{2})/a_{1}\lesssim\mathcal{O}(10)$ and relatively large octupole effects $0.01\lesssim\epsilon_{\rm oct}\lesssim 0.1$ (B. Liu et al., 2015). As a result of the strong octupole effects, mergers are possible across the full range of initial relative inclinations between the inner and orbital plane (right panel) and do not require the somewhat fine-tuned near-perpendicular configuration of the quadrupole-level approximation (e.g., needed for tertiary-driven binary black hole mergers which tend to have larger mass ratios, F. Antonini et al., 2017).
    Figure 7: Evolution of an example tripe that leads to an NSBH merger with large BH spin-orbit misalignment and significant residual eccentricity. The left panel shows large-amplitude ZKL oscillations of the NSBH eccentricity $e_{1}$ (black line) as a function of time since NSBH formation. Meanwhile, the direction of the NSBH orbital angular momentum oscillates while the BH spin direction remains effectively constant, which leads to large changes of the spin-orbit angle (red line). The system merges after a few cycles at around $t\approx 1.4\,\rm Myr$ with a spin-orbit misalignment close to $90\,\rm deg$. The right panel shows the eccentricity evolution as a function of the dominant GW frequency $f_{\rm GW}$. For most of the evolution it oscillates at low GW frequency potentially contributing to a stochastic background in the LISA bandwidth (Z. Xuan et al., 2024) until it reaches maximum eccentricity and undergoes an effectively isolated GW-driven inspiral.
    Figure 8: Same as Fig. 7 for an example system that merges before fully completing a full ZKL cycle. As a result the inner orbital angular momentum could only slightly deviate from the BH spin direction.
@@
    Non-adiabaticity causes large spin-orbit misalignments: The evolution of the BH spin can be characterised by the adiabaticity parameter $R=t_{\rm ZKL}/t_{\mathbf{S}_{\rm BH}}$ (K. R. Anderson et al., 2017; C. L. Rodriguez & F. Antonini, 2018) where $t_{\rm ZKL}$ is the ZKL timescale defined in Eq. (10) and
@@
    is the de Sitter precession timescale of the BH spin vector about the inner binary angular momentum. Since all NSBH mergers are driven by a nearby tertiary companion ($a_{2}(1-e_{2})/a_{1}\lesssim\mathcal{O}(10)$) and $R\propto\left(a_{2}\sqrt{1-e_{2}^{2}}\right)^{3}/a_{1}^{4}$ they occur in the non-adiabatic regime where $R\ll 1$ (left panel of Fig. 6). In this regime, the BH spin evolves too slowly to effectively follow the direction of the inner orbital angular momentum vector through the relativistic spin-orbit coupling. Thus, the BH spin direction remains essentially fixed whereas the inner binary orbital angular momentum vector evolves rapidly due to the ZKL oscillations caused by the nearby tertiary companion (C. L. Rodriguez & F. Antonini, 2018). As a result, the NSBHs can merge with a large spin-orbit misalignment of the BH, as exemplified by the system shown in Fig. 7 where—simultaneously to the NSBH eccentricity—the direction of its orbital angular momentum undergoes large-amplitude oscillation and eventually forms a large angle with the BH spin at merger. Nevertheless, closely aligned spin-orbit orientations can occur even in the non-adiabatic regime. This is the case if the tertiary perturbation is so strong that NSBHs are driven to merge before a full ZKL cycle could be completed and the orbital angular momentum could substantially depart from alignment, which is shown in an example in Fig. 8.
@@
    Tertiary-driven NSBH mergers can retain large residual eccentricity: In general, tertiary-driven NSBH mergers occur if the ZKL effect induces a critically large eccentricity ($e_{1}\approx 1)$ of the inner binary where the sudden energy-loss due to GW radiation at close periapsis shrinks the inner binary, decouples it from further perturbation from the tertiary, and is followed by an effectively isolated GW-driven inspiral (e.g., K. Silsbee & S. Tremaine, 2017). The critical eccentricity to undergo a GW-dominated inspiral can be either achieved before ZKL oscillations could be quenched by the relativistic apsidal precession of the inner binary or the latter becomes important before GW radiation dominates the binary evolution (C. L. Rodriguez & F. Antonini, 2018). In the former case the inner binary only undergoes a few ZKL cycles before it “plunges” directly into the highly eccentric GW-dominated regime, whereas in the latter case the binary is driven to a more “gentle” merger caused by the cumulative effect of damped eccentricity excitations across many ZKL cycles. As a result, the binary can retain a large eccentricity upon merging in the former case and typically circularises below measurability in the latter case. Which case applies can be estimated by comparing the angular momentum $j_{\rm GW}$ where GW radiation dominates over the ZKL effect with the value $j_{\rm GR}$ where the latter is quenched by the Schwarzschild precession of the inner binary, which are given by (F. Antonini et al., 2018; C. L. Rodriguez & F. Antonini, 2018)
@@
    Short delay times tie the tertiary-driven NSBH merger rate to the cosmic star formation: The median delay time from NSBH formation to merger ranges from $1.95\times 10^{4}$ (Bray) to $2.49\times 10^{5}\,\rm yr$ (Zero) and the maximum 95th percentile is found to be $2.58\times 10^{7}\,\rm yr$ (Zero). The short delay times can be understood by the large values of $j_{\rm GW}/j_{\rm GR}$ (see (iii)) which cause most mergers to occur in no more than a few octupole cycles $\sim\mathcal{O}(1)\times t_{\rm EKL}$ where $t_{\rm EKL}\sim t_{\rm ZKL}/\sqrt{\epsilon_{\rm oct}}$ is the timescale for octupole oscillations (often referred to as eccentric Kozai Lidov effect) (J. M. O. Antognini, 2015). In addition, massive stars evolve and form NSBHs after several Myr, which causes the total delay time from ZAMS to NSBH merger to be small on a cosmological scale and we expect the merger rate as a function of cosmological redshift to closely follow the cosmic star formation rate (P. Madau & M. Dickinson, 2014).
@@
    Tertiary-driven NSBH mergers form from close interacting binary stars: In order to survive the natal kick at NS formation the inner binaries need to be close. We find the inner semi-major axis distribution of surviving NSBH triples to peak at $a_{1}\sim\mathcal{O}(1)\,\rm AU$ and decline above a few $\sim 10\,\rm AU$ (see Fig. 2, red line). Thus, most of their progenitor stars underwent some sort of mass transfer (cf., blue patches) when they grew during stellar evolution and filled their Roche-lobe (P. P. Eggleton, 1983). Despite being driven by the gravitational perturbation from the tertiary companion, the properties of NSBH mergers (e.g., component masses and SN properties) are therefore strongly affected by the binary evolution of their progenitor stars (cf., P. Marchant & J. Bodensteiner, 2024; E. Laplace et al., 2025). This is different to the sample of NSBH mergers in triples obtained by G. Fragione & A. Loeb (2019a, b); A. A. Trani et al. (2022) who restricted their analyses to inner binaries with $a_{1}\gtrsim 10\,\rm AU$ that effectively evolve as single stars. Meanwhile, close and interacting inner binaries are more common (H. Sana et al., 2012, 2014), they are more likely to survive the NS natal kick, and the triple is less likely to be disrupted if the NS forms from a binary-stripped star, with reduced mass ejecta at SN explosion. Therefore, limiting the analyses to non-interacting triples overlooks most of the NSBH mergers.
@@
Figure 9: Distributions of the residual eccentricity at a dominant GW frequency $f_{\rm GW}=20\,\rm Hz$ (L. Wen, 2003) of NSBH mergers (left) and the final BH spin-orbit misalignment (right). Red patches show NSBH mergers in triples below a threshold of $e_{20}<0.4$. Tertiary-driven NSBH mergers with higher eccentricities could be observed as burst-like sources and are shown by yellow patches. The distribution of isolated NSBH mergers without tertiary companions is shown by blue patches. Orange, red, and teal lines show the cumulative distributions of all NSBH mergers in triples, the subset of those with $e_{20}<0.4$, and isolated binary mergers, respectively. The teal shaded region on the left panel show the detectability limit $e_{20}\gtrsim 0.07$ for LVK events.

### v2

@@ II.3 Gravitational three-body dynamics
Near a large eccentricity excitation of the inner binary the timescale for changes in the orbital elements could become shorter than the outer or inner orbital period, rendering the secular equations of motion inaccurate and likely underestimating the number of eccentric mergers by a few ten percent compared to direct three-body integration (e.g., Antonini et al., 2014; Grishin et al., 2018; Hamilton and Rafikov, 2024). Moreover, we do not attempt to model tidal deformation and mass-transfer of NSs which may alter the properties of NSBHs near the merger (Davies et al., 2005; Foucart et al., 2024; Zenati et al., 2025), which is beyond the scope of this work. Truncating the secular equations of motion at the octupole order we also neglect any higher-order terms which may become relevant for driving large eccentricities (Will, 2017). Following Will (2021) we consider the ratio between the timescales of octupole effects and the next-order hexadecapole terms
$\color[rgb]{0,0,0}\frac{T_{\rm oct}}{T_{\rm hex}}\sim\frac{a_{1}}{a_{2}e_{2}(1-e_{2}^{2})}\frac{1-3\eta}{\sqrt{1-4\eta}},$} (11)
where $\eta=m_{\rm NS}m_{\rm BH}/(m_{\rm NS}+m_{\rm BH})^{2}$. Since $m_{\rm NS}\ll m_{\rm BH}$ we find $(1-3\eta)/\sqrt{1-4\eta}\approx 1$. Thus, hexadecapole effects may become relevant for mildly hierarchical triples $a_{1}/a_{2}\lesssim 0.1$ and small outer eccentricities $e_{2}\lesssim 0.1$, which are suppressed by assuming a thermal distribution (see Sec. II.4).
We also evolve the sample of NSBH triples in one model (Maxwellian_100_10) with the direct $N$-body integrator MSTAR (Rantala et al., 2020; Mannerkoski et al., 2023), in order to estimate the impact of secular break-down at large eccentricities (Antonini et al., 2014) and truncation at octupole order on our results. MSTAR is based on an algorithmically regularized integration technique (e.g., Mikkola and Tanikawa, 1999; Preto and Tremaine, 1999; Mikkola and Merritt, 2008; Hellström and Mikkola, 2010; Trani and Spera, 2023) and uses the Gragg-Bulirsch-Stoer (GBS) extrapolation technique (Gragg, 1965; Bulirsch and Stoer, 1966) to evolve the equations of motion to 3.5 post-Newtonian order at high precision.
Using the output of any direct $N$-body integration to extract a physically meaningful eccentricity from the relative separation and velocity vectors $\mathbf{r}_{1}$ and $\mathbf{v}_{1}$, respectively, of merging compact object binaries is a non-trivial task. Obtaining the geometric eccentricity $e_{g}=(r_{\max}-r_{\min})/(r_{\max}+r_{\min})$ as utilised in eccentric waveform analyses (e.g., Shaikh et al., 2023) requires precise numerical methods to accurately resolve the apo- and pericentre $r_{\max}$ and $r_{\min}$, respectively, of the inpiralling binary. Meanwhile, calculating the Keplerian eccentricity form the Laplace–Runge–Lenz vector of the instantaneous vectors $\mathbf{r}_{1}$ and $\mathbf{v}_{1}$ (or post-Newtonian variants (Memmesheimer et al., 2004; Mannerkoski et al., 2019)), as widely done in previous population studies (e.g., Rodriguez et al., 2018; Zevin et al., 2021), can yield unphysically large eccentricity values if the binary begins to rapidly precess close to merger (Loutrel et al., 2019). Here, we do use the Keplerian eccentricity $e_{K}$ and semi-major axis $a_{K}$ calculated from
$\mathbf{e}_{K}$ $={\mathbf{v}_{1}\times(\mathbf{r}_{1}\times\mathbf{v}_{1})\over{G(m_{\rm BH}+m_{\rm NS})}}-{\mathbf{r}_{1}\over{{r}_{1}}}$ (12)
$\frac{1}{a_{K}}$ $=\frac{2}{r_{1}}-\frac{v_{1}^{2}}{G(m_{\rm BH}+m_{\rm NS})}.$ (13)
However, for this calculation we use the last $N$-body output ($\mathbf{r}_{1},\mathbf{v}_{1}$) of merging binaries where
$r_{1}$ $>200\frac{G(m_{\rm BH}+m_{\rm NS})}{c^{2}}{\quad\rm and}$ (14)
$e_{K}$ $>\mathcal{P}=\frac{G(m_{\rm BH}+m_{\rm NS})}{c^{2}}\frac{3-\eta}{a_{K}(1-e_{K}^{2})}.$ (15)
Crucially, requiring $e_{K}>\mathcal{P}$ in Eq. (15) corresponds to Will’s condition ensuring that the binary is still far enough from merging so that $e_{K}$ correctly traces circularisation upon gravitational-wave emission (Will, 2019), while Eq. (14) is imposed for numerical reasons as the direct $N$-body output of MSTAR can become sparse for separations closer to merger. We then proceed as Zevin et al. (2021) and use $e_{K}$ and $a_{K}$ as input to Peter’s equations (Peters, 1964) to obtain the eccentricity $e_{20}$ when the dominant gravitational-wave harmonics has a frequency of 20 Hz (Wen, 2003).
II.4 Initial conditions
@@
On the zero-age-main-sequence (ZAMS) we sample the initial mass of the primary star in the inner binary from an initial mass function $p(m_{1})\propto m_{1}^{-2.3}$ between $10$ and $100\,\rm M_{\odot}$ (Kroupa, 2001). We follow the observationally inferred distribution of massive binary stars by Sana et al. (2012) and sample the inner binary mass ratio $q_{1}=m_{2}/m_{1}$ from a power-law $p(q_{1})\propto q_{1}^{-0.1}$ and its orbital period from $p(\log P_{1}/{\rm days})\propto(\log P_{1}/{\rm days})^{-0.55}$. We reject any inner binary which would overflow its Roche-lobe at ZAMS. We assume that the inner binaries are initially circular ($e_{1}=0$) and sample the outer binary from a thermal eccentricity distribution, a log-uniform semi-major axis distribution, and the outer binary mass ratio $q_{2}=m_{3}/m_{1}$ from a broken power-law $p(q_{2})\propto q_{2}^{-1.2}$ below $q_{2}<0.3$ and $\propto q_{2}^{-2.0}$ above $q_{2}<0.3$ (Moe and Di Stefano, 2017), limited to $0.08\,{\rm M_{\odot}}<m_{3}<5\,\rm{\rm M_{\odot}}$. For any given inner binary, we repeat sampling the outer binary until the triple is dynamically stable (Mardling and Aarseth, 2001). We then discard any triple with an outer semi-major axis above $a_{2}>5\times 10^{4}\,\rm AU$ because they are likely disrupted by stellar fly-bys and tides from the host galaxy (Jiang and Tremaine, 2010; El-Badry and Rix, 2018; Hamilton and Modak, 2024; Stegmann et al., 2024). Finally, the orbital orientations of the inner and outer orbital frame are sampled from random isotropic distributions, leading to a random mutual inclination between the inner and outer orbit. Using the procedure described above, we construct a total number of $N_{\rm tot}=10^{7}$ stellar triples which we evolve in this work.
@@ III Results
Starting on the ZAMS, the orbits of stellar triples have to survive two consecutive SNe and any inner binary mass transfer episode in order to successfully develop stable NSBH triples that could eventually lead to tertiary-driven mergers. In Fig. 2 we show how the properties of surviving systems change across different times during the stellar evolution. We find that a few $\sim 0.1\,\%$ of all stellar triples initiated at ZAMS (grey patches) survive the stellar evolution and form a stable NSBH triple (red lines) and about $\sim 40\,\%$ thereof subsequently lead to a tertiary-driven merger (red patches). The latter fraction is similar to the merger fraction found by Fragione and Loeb (2019b) for their population of wider NSBH triples. During stellar evolution the NS formation at second SN is the most detrimental event to the triple stability and critically truncates the parameter distribution of the survivors. On the one hand, the natal kick of the newly formed NS poses the risk to disrupt the inner binary producing a dearth of stable NSBH inner binaries (blue line) above $a_{1}\gtrsim 10^{2}\,\rm AU$ (left panel). On the other hand, the mass-loss associated with the SN imparts a kick to the inner binary centre-of-mass which causes a steep decline in the distribution of $a_{2}/a_{1}$ of surviving NSBH triples above $\gtrsim 10^{2}$ (right panel). The stellar evolutionary constraints are further illustrated in Fig. 3 showing that the surviving NSBH triples occupy a somewhat narrow strip in $a_{2}$ versus $a_{1}$ which is predominantly determined by the mass-loss associated with the second SN and the need for the outer orbit to survive it. Consequently, we find different natal kick models only to affect the upper limit on the inner binary semi-major axis of surviving NSBH triples but not substantially changing the semi-major axis ratio distribution, as shown in Fig. 9.
@@
As a result, the surviving NSBH triples naturally populate a distinct region of parameter space that facilitates the formation of tertiary-driven mergers with spin-orbit misalignment and residual eccentricity. On the one hand, the SN kick imparted to the newly formed NS disrupts triples in wide architectures and is only withstood by triples with relatively small inner and outer semi-major axes. One the other hand, NSBH triples are characterised by a small inner binary mass ratio $q=m_{\rm NS}/m_{\rm BH}\ll 1$ since $m_{\rm NS}\ll m_{\rm BH}$. Therefore, we obtain tertiary-driven NSBH mergers that are, by and large, characterised by (i) strong octupole effects, (ii) a non-adiabatic BH spin evolution, (iii) a peculiar dynamics in which the NSBH mergers “plunge” (Rodriguez and Antonini, 2018) directly into a highly eccentric, GW-dominated phase without being significantly quenched by apsidal Schwarzschild precession of the inner binary, and (iv) short delay times from ZAMS to merger. Moreover, (v) the evolution of the NSBH progenitor stars is dominated by close stellar binary interactions, as explained below.
@@
    Octupole effects enable mergers across a wide range of initial relative inclinations: The secular equations of motion of hierarchical triples are obtained by expanding the three-body Hamiltonian in semi-major axis ratio $a_{1}/a_{2}$ (e.g., Harrington, 1968). To the lowest (“quadrupole”) order large inner binary eccentricities are only excited if the relative inclination between the inner and outer orbital planes is near-perpendicular. Large eccentricities across a much wider range of relative inclinations can be excited if the next-order (“octupole”) terms becomes relevant (e.g., Ford et al., 2000; Liu et al., 2015; Naoz, 2016) whose relative strength can be characterised as (Naoz et al., 2013)
@@
    As shown in Fig. 4, $\epsilon_{\rm oct}$ of the surviving triples can become as large as $\epsilon_{\rm oct}\lesssim 0.1$ (left panel) because the inner binary masses of NSBHs are very unequal, the semi-major axis ratio of surviving NSBH triples is relatively compact, and the thermal eccentricity distribution of the outer binary disfavours circular outer orbits (middle panel). NSBH mergers predominantly occur with nearby tertiaries where $a_{2}(1-e_{2})/a_{1}\lesssim\mathcal{O}(10)$ and relatively large octupole effects $0.01\lesssim\epsilon_{\rm oct}\lesssim 0.1$ (Liu et al., 2015). As a result of the strong octupole effects, mergers are possible across the full range of initial relative inclinations between the inner and orbital plane (right panel) and do not require the somewhat fine-tuned near-perpendicular configuration of the quadrupole-level approximation (e.g., needed for tertiary-driven binary black hole mergers which tend to have larger mass ratios, Antonini et al., 2017).
    Figure 6: Evolution of an example tripe that leads to an NSBH merger with large BH spin-orbit misalignment and significant residual eccentricity. The left panel shows large-amplitude ZKL oscillations of the NSBH eccentricity $e_{1}$ (black line) as a function of time since NSBH formation. Meanwhile, the direction of the NSBH orbital angular momentum oscillates while the BH spin direction remains effectively constant, which leads to large changes of the spin-orbit angle (red line). The system merges after a few cycles at around $t\approx 0.2\,\rm Myr$ with a spin-orbit misalignment close to $90\,\rm deg$. The right panel shows the eccentricity evolution as a function of the dominant GW frequency $f_{\rm GW}$. Initial parameters of the system are $m_{1}=11.14\,\rm M_{\odot}$, $m_{2}=1.50\,\rm M_{\odot}$, $m_{3}=0.21\,\rm M_{\odot}$, $e_{1}=0.97$, $a_{1}=2.99\,\rm AU$, $e_{2}=0.19$, $a_{2}=27.57\,\rm AU$, $\mathbf{\hat{e}}_{1}=(0.36,0.93,-0.11)^{\rm T}$, $\mathbf{\hat{j}}_{1}=(-0.53,0.10,-0.84)^{\rm T}$, $\mathbf{\hat{e}}_{2}=(-0.57,-0.39,0.72)^{\rm T}$, $\mathbf{\hat{j}}_{2}=(0.68,0.28,0.68)^{\rm T}$, and $\mathbf{\hat{S}}_{\rm BH}=(-0.70,0.52,-0.49)^{\rm T}$.
@@
    Non-adiabaticity causes large spin-orbit misalignments: The evolution of the BH spin can be characterised by the adiabaticity parameter $R=t_{\rm ZKL}/t_{\mathbf{S}_{\rm BH}}$ (Anderson et al., 2017; Rodriguez and Antonini, 2018) where $t_{\rm ZKL}$ is the ZKL timescale defined in Eq. (10) and
@@
    is the de Sitter precession timescale of the BH spin vector about the inner binary angular momentum. Since all NSBH mergers are driven by a nearby tertiary companion ($a_{2}(1-e_{2})/a_{1}\lesssim\mathcal{O}(10)$) and $R\propto\left(a_{2}\sqrt{1-e_{2}^{2}}\right)^{3}/a_{1}^{4}$ they occur in the non-adiabatic regime where $R\ll 1$ (left panel of Fig. 5). In this regime, the BH spin evolves too slowly to effectively follow the direction of the inner orbital angular momentum vector through the relativistic spin-orbit coupling. Thus, the BH spin direction remains essentially fixed whereas the inner binary orbital angular momentum vector evolves rapidly due to the ZKL oscillations caused by the nearby tertiary companion (Rodriguez and Antonini, 2018). As a result, the NSBHs can merge with a large spin-orbit misalignment of the BH, as exemplified by the system shown in Fig. 6 where—simultaneously to the NSBH eccentricity—the direction of its orbital angular momentum undergoes large-amplitude oscillation and eventually forms a large angle with the BH spin at merger. Nevertheless, closely aligned spin-orbit orientations can occur even in the non-adiabatic regime. This is the case if the tertiary perturbation is so strong that NSBHs are driven to merge before a full ZKL cycle could be completed and the orbital angular momentum could substantially depart from alignment, which is shown in an example in Fig. 10.
    Exemplified in Figs. 6 and 10 we also note that for most of the evolution the binaries oscillate at low GW frequency potentially being resolvable or contributing to a stochastic background in the LISA bandwidth (Xuan et al., 2024; Knee et al., 2024; Xuan et al., 2025) until it reaches maximum eccentricity and undergoes an effectively isolated GW-driven inspiral.
@@
    Tertiary-driven NSBH mergers can retain large residual eccentricity: In general, tertiary-driven NSBH mergers occur if the ZKL effect induces a critically large eccentricity ($e_{1}\approx 1)$ of the inner binary where the sudden energy-loss due to GW radiation at close periapsis shrinks the inner binary, decouples it from further perturbation from the tertiary, and is followed by an effectively isolated GW-driven inspiral (e.g., Silsbee and Tremaine, 2017). The critical eccentricity to undergo a GW-dominated inspiral can be either achieved before ZKL oscillations could be quenched by the relativistic apsidal precession of the inner binary or the latter becomes important before GW radiation dominates the binary evolution (Rodriguez and Antonini, 2018). In the former case the inner binary only undergoes a few ZKL cycles before it “plunges” directly into the highly eccentric GW-dominated regime, whereas in the latter case the binary is driven to a more “gentle” merger caused by the cumulative effect of damped eccentricity excitations across many ZKL cycles. As a result, the binary can retain a large eccentricity upon merging in the former case and typically circularises below measurability in the latter case. Which case applies can be estimated by comparing the angular momentum $j_{\rm GW}$ where GW radiation dominates over the ZKL effect with the value $j_{\rm GR}$ where the latter is quenched by the Schwarzschild precession of the inner binary, which are given by (Antonini et al., 2018; Rodriguez and Antonini, 2018)
@@
    Short delay times tie the tertiary-driven NSBH merger rate to the cosmic star formation: The median delay time from NSBH formation to merger ranges from $1.95\times 10^{4}$ (Bray) to $2.49\times 10^{5}\,\rm yr$ (Zero) and the maximum 95th percentile is found to be $2.58\times 10^{7}\,\rm yr$ (Zero). The short delay times can be understood by the large values of $j_{\rm GW}/j_{\rm GR}$ (see (iii)) which cause most mergers to occur in no more than a few octupole cycles $\sim\mathcal{O}(1)\times t_{\rm EKL}$ where $t_{\rm EKL}\sim t_{\rm ZKL}/\sqrt{\epsilon_{\rm oct}}$ is the timescale for octupole oscillations (often referred to as eccentric Kozai Lidov effect) (Antognini, 2015). In addition, massive stars evolve and form NSBHs after several Myr, which causes the total delay time from ZAMS to NSBH merger to be small on a cosmological scale and we expect the merger rate as a function of cosmological redshift to closely follow the cosmic star formation rate (Madau and Dickinson, 2014).
@@
    Tertiary-driven NSBH mergers form from close interacting binary stars: In order to survive the natal kick at NS formation the inner binaries need to be close. We find the inner semi-major axis distribution of surviving NSBH triples to peak at $a_{1}\sim\mathcal{O}(1)\,\rm AU$ and decline above a few $\sim 10\,\rm AU$ (see Fig. 2, red line). Thus, most of their progenitor stars underwent some sort of mass transfer (cf., blue patches) when they grew during stellar evolution and filled their Roche-lobe (Eggleton, 1983). Despite being driven by the gravitational perturbation from the tertiary companion, the properties of NSBH mergers (e.g., component masses and SN properties) are therefore strongly affected by the binary evolution of their progenitor stars (cf., Marchant and Bodensteiner, 2024; Laplace et al., 2025). This is different to the sample of NSBH mergers in triples obtained by Fragione and Loeb (2019b, a); Trani et al. (2022) who restricted their analyses to inner binaries with $a_{1}\gtrsim 10\,\rm AU$ that effectively evolve as single stars. Meanwhile, close and interacting inner binaries are more common (Sana et al., 2012, 2014), they are more likely to survive the NS natal kick, and the triple is less likely to be disrupted if the NS forms from a binary-stripped star, with reduced mass ejecta at SN explosion. Therefore, limiting the analyses to non-interacting triples overlooks most of the NSBH mergers.
@@
Figure 7: Distributions of the residual eccentricity at a dominant GW frequency $f_{\rm GW}=20\,\rm Hz$ (Wen, 2003) of NSBH mergers (left) and the final BH spin-orbit misalignment (right). Red patches show NSBH mergers in triples below a threshold of $e_{20}<0.4$. Tertiary-driven NSBH mergers with higher eccentricities could be observed as burst-like sources and are shown by yellow patches. The distribution of isolated NSBH mergers without tertiary companions is shown by blue patches. Orange, red, and teal lines show the cumulative distributions of all NSBH mergers in triples, the subset of those with $e_{20}<0.4$, and isolated binary mergers, respectively. Dashed lines show the cumulative eccentricity distribution from our direct $N$-body integration (see Sec. II.3). The teal shaded region on the left panel show the detectability limit $e_{20}\gtrsim 0.07$ for LVK events.